• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 11:18:09 on October21,2009

    Entry number 295966

    keyword=Compton laser update

    Compton laser can be turned on now. It looks like we have to wait for some time for the laser to be warmed up (according to Mark's axperience may be couple of hours). Right now laser spot is moving in the quadrant detectors, we will check if it stop moving and then try to lock it. Hopefully there is not much difficulty with locking.